A boiler takes in feed water at 70 °C and generates steam at a pressure of 10 bar and dryness fraction of 0.92. The wet steam enters a superheater and emerges with the same pressure at a temperature of 400 °C. If the steam flow rate is 5 tonnes per hour, determine (i) the rate at which heat transfer takes place in the boiler and (ii) the rate at which heat transfer takes place in the superheater.
1
Expert's answer
2018-08-14T10:17:08-0400
i)
h_1=293.1 kJ/kg
h_2=2616.9 kJ/kg
P_1=(h_2-h_1 ) dm/dt= (2616.9-293.1)(5000/3600)=3227.5 kW.
ii)
h_3=3262.8 kJ/kg
P_2=(h_3-h_2 ) dm/dt= (3262.8-2616.9)(5000/3600)=897 kW.